# Coupled Laplacian Eigenmaps for Locally-Aware 3D Rigid Point Cloud Matching

### v1

@@ Abstract
Point cloud matching, a crucial technique in computer vision, medical and robotics fields, is primarily concerned with finding correspondences between pairs of point clouds or voxels. In some practical scenarios, emphasizing local differences is crucial for accurately identifying a correct match, thereby enhancing the overall robustness and reliability of the matching process. Commonly used shape descriptors have several limitations and often fail to provide meaningful local insights on the paired geometries. In this work, we propose a new technique, based on graph Laplacian eigenmaps, to match point clouds by taking into account fine local structures. To deal with the order and sign ambiguity of Laplacian eigenmaps, we introduce a new operator, called Coupled Laplacian, that allows to easily generate aligned eigenspaces for multiple rigidly-registered geometries. We show that the similarity between those aligned high-dimensional spaces provides a locally meaningful score to match shapes. We initially evaluate the performance of the proposed technique in a point-wise manner, specifically focusing on the task of object anomaly localization using the MVTec 3D-AD dataset. Additionally, we define a new medical task, called automatic Bone Side Estimation (BSE), which we address through a global similarity score derived from coupled eigenspaces. In order to test it, we propose a benchmark collecting bone surface structures from various public datasets. Our matching technique, based on Coupled Laplacian, outperforms other methods by reaching an impressive accuracy on both tasks. The code to reproduce our experiments is publicly available at https://github.com/matteo-bastico/CoupledLaplacian and in the Supplementary Code.
@@ 1 Introduction
Over the years, several methods have been proposed to tackle the challenge of accurate and efficient 3D shape matching and retrieval [9, 5, 11, 53, 51, 65, 73]. Data-driven 3D shape descriptors [54], capturing underlying properties of the shapes under study, are the common denominator of early shape matching techniques. Global descriptors, such as volume and areas descriptors [78], describe the entirety of the shape in one compact representation, often failing to capture local fine details of complex geometries. On the other hand, local descriptors [39, 57] aim to tackle this issue but they generally are sensitive to noise, based on landmarks and they might not capture semantic information [63]. More recently, deep-learned shape descriptors [72, 6] and deep neural networks for shape matching, based on auto-encoders [73] or transformers [59, 68], have been proposed. Despite their good performances, these methods require a huge amount of annotated data for training, which are hard to collect in fields such as medical imaging [38]. Furthermore, non-rigid point cloud matching and retrieval methods [36, 37, 71] are designed to handle shape deformations and, therefore, they might be excessively flexible ignoring fine local details that are not due to deformations, such as anomalies.
@@
In this study, we introduce a novel method for rigid 3D point cloud matching, based on spectral Laplacian eigenmaps [4, 24], which focus on local details. This technique overcomes the previously mentioned limitations and constraints associated to global and local descriptors and, although it can be seen as a linear graph neural network, it does not require training. One of the main reasons that combinatorial and geometric Laplacians are often considered for spectral graph processing is that their eigenvectors possess properties like the classical Fourier basis functions [79]. We exploit these properties to perform locally-aware shape comparison of point clouds, equipped with k-nearest neighbor graph, without relying on additional descriptors. Nevertheless, eigenspaces alignment, including both eigenvalue ordering and sign disambiguity of the eigendecomposition, is required to correctly match shapes in the spectral space. Current State-of-The-Art methods for such alignment are based on Laplacian eigenfunctions matching [42, 60], i.e. matching the histograms of the eigenvectors. The latter is not robust and frequently fails for highly symmetric geometries, such as bones, with small eigenvalue separation. In this work, we introduce a novel operator, called Coupled Laplacian, designed to simultaneously produce aligned eigenspaces for multiple registered geometries. Namely, we show that when two or more shapes, each with their corresponding graphs, are merged into a single graph using artificial cross-edges, the eigendecompositon of the Laplacian derived from such a combined graph yields aligned spectral spaces for each individual component. Furthermore, this method is order-invariant, count-agnostic and landmarks-free, enabling it to handle point clouds without being influenced by their specific arrangement or the number of points contained. Finally, we utilize the distance between these aligned higher-dimensional spectral spaces, which accounts for intricate local structures, as a global or point-wise score for different applications of 3D shape matching, as shown in Fig. 1.
@@
Hence, the identification of local differences between two shapes, using the proposed matching technique, allows us to perform 3D object anomaly detection in the spectral domain with high accuracy. We tested the performances of our method on the MVTec 3D-AD dataset [7], recently proposed for unsupervised point cloud industrial anomaly detection and localization. Furthermore, in the context of medical imaging, correctly identifying the side of a bone (left or right side) is crucial for diagnosis, treatment planning, or skeletal analysis [12]. We refer to the task of determining the side of a Bone or anatomical structure as laterality or Side Estimation (BSE). Recent studies on bone landmark detection [19] and statistical shape modeling [32] have highlighted the limitations of manual side identification for ensuring proper functioning. To the best of our knowledge, manual bone markings [2] are currently the only technique used for this purpose. Consequently, automatic laterality determination arises as an interesting, yet challenging, task to assist the development of fully automated pipelines for medical image analysis, such as patient-specific preoperative planning for Total Knee Arthroplasty (TKA) [34] or Anterior Cruciate Ligament Reconstruction (ACLR) [18, 44]. Thanks to the bilateral symmetry of the animal and human body [28, 67], i.e. right and left sides are mirror shapes of one another, the BSE can be defined as a non-trivial shape matching problem. Given a reference, or source, of known side, to assess the side of a target bone, it is enough to find the best match between the source-target pair and its mirrored counterpart, or vice versa. The complexity of this task is a consequence of subtle local differences present in mirrored bones. This makes it an ideal application of the proposed shape matching technique that utilizes Coupled Laplacian eigenmaps as local geometry descriptors. We generate a benchmark for human BSE by extracting surface point clouds of different bones, e.g. femur, hip and tibia, from public datasets [19, 32, 47]. We also discuss, in the Supplementary Material, the cross-species generalization of the proposed matching algorithm from human to animal, and vice-versa, on an internal dataset [40].
@@ 2 Related Works
Early efforts in developing hand-crafted 3D local features for point cloud matching, registration, recognition and retrieval have typically drawn inspiration from 2D descriptors. Several methods, such as Signature of Histograms of OrienTations (SHOT) [58, 66], Rotational Projection Statistics (RoSP) [26] and Unique Shape Context (USC) [66] rely on the estimation of an unique Local Reference Frame (LRF). The latter is usually based on the eigendecomposition of the covariance matrix of the neighbours of a point of interest, which are then projected into the LRF to analyze their geometric properties. For instance, SHOT [58] captures local shape and orientation information by computing histograms of surface normals and point distribution around keypoints. In contrast, LRF-free approaches [10, 56, 57] try to rely just on features that are intrinsically invariant. The most common LRF-free features are Fast Point Feature Hisograms (FPFH) [57] which, similarly to SHOT, calculates histograms of surface normals considering their relationships in a local region around a keypoint. Despite the progress made with hand-crafted 3D local features, they encounter difficulties when it comes to deal with issues like point cloud resolutions, noisy data, and occlusions [27].
@@
On the contrary, spectral-based point cloud descriptors [1, 29, 52, 70, 62, 55, 69] are a category of feature extraction methods that leverage spectral analysis techniques from graph theory and linear algebra to capture the underlying structure and intrinsic geometric properties of point clouds. Among them, shape-DNA [52] is a surface and solid descriptor based on eigenvalue analysis of the Laplace-Beltrami operator. They propose to use the sequence of eigenvalues (spectrum) of the Laplace operator as a fingerprint characterizing the intrinsic geometry of 3D shapes represented as point clouds. This method has been used for shape retrieval, classification, and correspondence. Weinmann et al. [70] also proposed to extract a features set consisting of $8$ eigenvalues-based indices for each 3D point of a cloud. Furthermore, Heat Kernel Signature (HKS) [62] and Wave Kernel Signatures (WKS) [1] are spectral-based descriptors measuring how heat and wave propagate across a shape, having the eigendecomposition as leading element of the computation. Scaled eigenvectors evaluated at each point are instead directly exploited by the Global Point Signature (GPS) [55] to represent a point cloud as a set of infinite-dimensional vectors, characterizing each point within the global context of the surface it belongs. Nevertheless, the vast majority of these works assumes that the eigenvalues of a shape are distinct and, therefore, can be ordered. Indeed, in practice, due to numerical approximations, we cannot guarantee that the eigenvalues of the Laplacian are all distinct and, possible symmetries in the shapes may cause some of them to have multiplicity grater than one. As shown by Mateus et al. [42], when dealing with shape matching, an elegant way to overcome this problem is to use the Laplacian eigenmaps scheme [4], which can be seen as a reduced GPS, and perform a-posteriori alignment of the resulting point cloud embeddings. Matching the eigenfunctions histograms, i.e. their signatures, is the only reliable method for such embeddings alignment [42, 60]. Unfortunately, when the eigenvalues separation is small due to symmetries in the geometries, this method becomes unstable and sensitive to noise. To overcome this issue, in this work we propose the Coupled Laplacian operator to produce a-priori aligned eigenmaps for several registered similar shapes, and perform locally-aware shape matching.
@@
More recently, deep-learning based point-cloud descriptors [25, 72, 6, 77] arose as an alternative to generate local features for 3D sufraces. Here we can distinguish three main categories depending on the backbone architecture they employ. Convolutional Neural Networks (CNNs) are often used on point clouds projected into 2D depth images [17, 30] or directly on 3D voxels [25, 77, 30]. For example, 3DMatch [77] aims to learn a local volumetric patch descriptor for establishing correspondences between partial 3D data. Secondly, to work directly on raw point cloud data, PointNet and PointNet++ have been proposed aiming to learn rotation and permutation invariant features [49, 50]. Based on PointNet, several learned descriptors have been proposed [15, 14, 75, 76]. Among them, PPFNet [15] and PPF-FoldNet [14] try to improve the feature representation of PointNet by incorporating global context and point-pair features. Nevertheless, the lack of convolutional layers in these models limits the learning of local geometries. Finally, transformers-based descriptors, such as Deep Closest Point [68], have been recently proposed trying to exploit the attention mechanisms to capture shapes and surfaces intrinsic characteristics. Arguably, training a deep-learning model, especially if based on transformers, it is not always feasible in terms of training samples required to achieve good performances. Furthermore, in supervised algorithms [77, 25, 15, 14], paired 3D patches, such as in the 3DMatch dataset [77], are needed to train the models. For this reason, the 3D rigid point cloud matching method we propose here does not need training and utilizes only the similarity among properly aligned spectral embeddings to relate an unseen target shape to the given references.
@@ 3 Method
where $\boldsymbol{B}\in\mathbb{R}^{n\times n}$ is generally set as $\boldsymbol{D}$, and $\lambda_{i}\leq\lambda_{i+1}$. Finally, eigenmaps are simply eigenvectors sub-spaces, generated by leaving out $\boldsymbol{\phi}_{0}$, corresponding to $\lambda_{0}=0$ and using the next $m$ eigenvectors for embedding graph nodes in an $m$-dimensional space, $\boldsymbol{x}_{i}\rightarrow[\boldsymbol{\phi}_{1}(i),\cdots,\boldsymbol{\phi}_{m}(i)]$
@@
Coupled Laplacian. The complete pipeline to generate aligned Laplacian embeddings through graph coupling is illustrated in Fig. 2. Let $\mathcal{G}^{\mathcal{T}}$ be the graph corresponding to a target 3D point cloud and $\mathcal{G}^{S_{1}},\cdots,\mathcal{G}^{S_{N}}$ the ones obtained from $N$ different sources. We construct a global graph, or coupled graph, $\mathcal{G}^{C}$, by adding cross-connections separately between the vertices of the source, $\mathcal{V}^{\mathcal{T}}=\{\boldsymbol{x}^{\mathcal{T}}_{i}\}_{i=1}^{n^{\mathcal{T}}}$, and each of the reference shapes, $\mathcal{V}^{S_{k}}=\{\boldsymbol{x}^{S_{k}}_{i}\}_{i=1}^{n^{S_{k}}}$ with $k=1,\cdots,N$. To include meaningful cross-connections, we first perform a rigid [20, 21, 45] or affine [45] registration of the source geometries to better match the target. The latters are preferred to a non-rigid registration because, when dealing with the identification of fine local variations, they do not change the relative position of the points inside a point-set and, therefore, the deformations are kept almost unchanged. After that, a sub-set of vertices, $F^{\mathcal{T}}\subset\mathcal{V}^{\mathcal{T}}$ , is stochastically extracted from the target geometry and its nearest correspondence is searched in each of the aligned reference point clouds, such as
$F^{S_{k}}=\{\boldsymbol{f}^{S_{k}}_{j}:\min_{\boldsymbol{x}^{S_{k}}_{i}\in\mathcal{V}^{S_{k}}}\text{d}^{2}(\boldsymbol{x}^{S_{k}}_{i},\boldsymbol{f}^{\mathcal{T}}_{j})|\boldsymbol{f}^{\mathcal{T}}_{j}\in F^{\mathcal{T}}\}$ (3)
@@
with $k=1,\cdots,N$. In this way, there is no constraint on the original number points comprising each shape nor on their initial coordinates system, e.g. camera coordinates. Furthermore, the cardinality of the target sub-set, is chosen as a fraction of its total points, $|F^{\mathcal{T}}|=l\cdot n^{\mathcal{T}}$, where $0<l\leq 1$. The cross-connection are therefore added for each pair $(\boldsymbol{f}^{S_{k}}_{i},\boldsymbol{f}^{\mathcal{T}}_{i})$, $1\leq i\leq|F^{\mathcal{T}}|$, between the target and the $k$-th source. Finally, we arrange the vertex indices of the coupled graph such that they are grouped for each individual geometry, and the whole weighted adjacency matrix can be computed as in Eq. 1.
@@
We define the Laplacian matrix of a global graph, constructed as describe above, as the Coupled Laplacian, $\boldsymbol{L}^{C}\in\mathbb{R}^{n\times n}$, where $n=n^{\mathcal{T}}+\sum_{k=1}^{N}n^{S_{k}}$. Thanks to the vertex ordering of the coupled graph, the solution of the generalized eigenproblem of Eq. 2 applied to the Coupled Laplacian with $\boldsymbol{B}^{C}=\text{diag}([\boldsymbol{D}^{\mathcal{T}},\boldsymbol{D}^{S_{1}},\cdots,\boldsymbol{D}^{S_{N}}])$, which is the stack of uncoupled diagonal degree matrices, gives coupled eigenvectors $\{\boldsymbol{\phi}^{C}_{i}\}_{i=0}^{n}$. These eigenvectors can then be split into each single component of the coupled graph as
@@
Shape Matching with Eigenmaps. The Coupled Laplacian allows the generation of $m$-dimensional embeddings aligned for the graph vertices of the target and reference shapes, $\boldsymbol{\Phi}^{\mathcal{T}}=(\boldsymbol{\phi}^{\mathcal{T}}_{1},\cdots,\boldsymbol{\phi}^{\mathcal{T}}_{m})\in\mathbb{R}^{n_{T}\times m}$ and $\boldsymbol{\Phi}^{S_{k}}=(\boldsymbol{\phi}^{S_{k}}_{1},\cdots,\boldsymbol{\phi}^{S_{k}}_{m})\in\mathbb{R}^{n_{S_{k}}\times m}$ for $1\leq k\leq N$, respectively. Therefore, a comparison of multiple geometries in this higher dimensional spectral space yields to a proper consideration of local structures. We consider in the following two possible applications of the coupled eigenmaps: (1) a global shape matching score and (2) local similarity scores. The first one can be obtained by measuring the similarity between the aligned eigenspaces through the Grassmann distance, $d_{G}(\cdot,\cdot)$ [74]. Nevertheless, it cannot be computed directly, since we consider point-sets of arbitrary size and we do not have point-to-point correspondences. Hence, we restrict the distance computation only to the set of cross-connected vertices, $F^{\mathcal{T}}$ for the target and $F^{S_{k}}$ for the $k$-th reference. In order to get the reduced basis restricted only to those coupled points, we perform a QR decomposition of the restriction of the eigenmodes to the cross-points. The best matching source is therefore given by
@@
where $\boldsymbol{Q}^{\mathcal{T}}$ and $\boldsymbol{Q}^{S_{k}}$ are obtained from the QR factorization of $\boldsymbol{\Phi}^{\mathcal{T}}(F^{\mathcal{T}},:)$ and $\boldsymbol{\Phi}^{S_{k}}(F^{S_{k}},:)$, respectively. This first approach is used in the following to solve the BSE task by using two source shapes, i.e. the reference bone of known side and its mirrored version.

### v2

@@ Abstract
Point cloud matching, a crucial technique in computer vision, medical and robotics fields, is primarily concerned with finding correspondences between pairs of point clouds or voxels. In some practical scenarios, emphasizing local differences is crucial for accurately identifying a correct match, thereby enhancing the overall robustness and reliability of the matching process. Commonly used shape descriptors have several limitations and often fail to provide meaningful local insights about the paired geometries. In this work, we propose a new technique, based on graph Laplacian eigenmaps, to match point clouds by taking into account fine local structures. To deal with the order and sign ambiguity of Laplacian eigenmaps, we introduce a new operator, called Coupled Laplacian [^1] Code: https://github.com/matteo-bastico/CoupLap, that allows to easily generate aligned eigenspaces for multiple registered geometries. We show that the similarity between those aligned high-dimensional spaces provides a locally meaningful score to match shapes. We firstly evaluate the performance of the proposed technique in a point-wise manner, focusing on the task of object anomaly localization on the MVTec 3D-AD dataset. Additionally, we define a new medical task, called automatic Bone Side Estimation (BSE), which we address through a global similarity score derived from coupled eigenspaces. In order to test it, we propose a benchmark collecting bone surface structures from various public datasets. Our matching technique, based on Coupled Laplacian, outperforms other methods by reaching an impressive accuracy on both tasks.
footnotetext: Corresponding author: matteo.bastico@minesparis.psl.eu
@@ 1 Introduction
Over the years, several methods have been proposed to tackle the challenge of accurate and efficient 3D shape matching and retrieval [9, 5, 11, 53, 51, 65, 73]. Data-driven 3D shape descriptors [54], capturing underlying properties of the shapes under study, are the common denominator of early shape matching techniques. Global descriptors, such as volume and areas descriptors [78], describe the entirety of the shape in one compact representation, often failing to capture local fine details of complex geometries. On the other hand, local descriptors [38, 57] aim to tackle this issue but they generally are sensitive to noise, based on landmarks and they might not capture semantic information [63]. More recently, deep-learned shape descriptors [72, 6] and neural networks for shape matching, based on auto-encoders [73] or transformers [59, 68], have been proposed. Despite their good performances, these methods require a huge amount of annotated data for training, which are hard to collect in fields such as medical imaging [37]. Furthermore, non-rigid point cloud matching and retrieval methods [35, 36, 71] are designed to handle shape deformations and, therefore, they might be excessively flexible ignoring fine local details that are not due to deformations, such as anomalies.
@@
In this study, we introduce a novel method for rigid 3D point cloud matching, based on spectral Laplacian eigenmaps [4, 23], which focus on local details. This technique is designed to overcome the limitations of shape descriptors and, although it can be seen as a linear graph neural network, it does not require training. One of the main reasons that combinatorial and geometric Laplacians are often considered for spectral graph processing is that their eigenvectors possess properties like the classical Fourier basis functions [79]. We leverage these characteristics to perform locally-aware shape comparison of point clouds, equipped with k-nearest neighbor graph, without relying on additional descriptors. Nevertheless, eigenspaces alignment, including both eigenvalue ordering and sign disambiguity of the eigendecomposition, is required to correctly match shapes in the spectral space. Current State-of-The-Art methods for such alignment are based on Laplacian eigenfunctions matching [42, 60], i.e. matching the histograms of the eigenvectors. The latter is not robust and frequently fails in possible scenarios with highly symmetric geometries, such as bones, with small eigenvalue separation. In this work, we introduce a novel operator, called Coupled Laplacian, designed to simultaneously produce aligned eigenspaces for multiple registered geometries. Namely, we show that when two or more shapes, each with their corresponding graphs, are merged into a single graph using artificial cross-edges, the eigendecompositon of the Laplacian derived from such a combined graph yields aligned spectral spaces for each individual component. Furthermore, this method is order-invariant, count-agnostic and landmarks-free, enabling it to handle point clouds without being influenced by their specific arrangement or the number of points contained. Finally, we utilize the distance between these aligned higher-dimensional spectral spaces, which accounts for intricate local structures, as a global or point-wise score for different applications of 3D shape matching, as shown in Fig. 1.
@@
One natural application of the proposed technique, is 3D object anomaly detection thought the identification of local differences between two shapes. Hence, we tested our method on the MVTec 3D-AD dataset [7], recently proposed for unsupervised point cloud anomaly detection and localization. Furthermore, in the context of medical imaging, correctly identifying the side of a bone (left or right) is crucial for diagnosis, treatment planning, or skeletal analysis [12]. We refer to this task as Bone Side Estimation (BSE). Recent studies on bone landmark detection [18] and statistical shape modeling [31] have highlighted the limitations of manual side identification for ensuring proper functioning. To the best of our knowledge, manual bone markings [2] are currently the only technique used for this purpose. Consequently, automatic BSE arises as an interesting, yet challenging, task to assist the development of fully automated pipelines for medical image analysis, such as patient-specific preoperative planning for Total Knee Arthroplasty (TKA) [33] or Anterior Cruciate Ligament Reconstruction (ACLR) [17, 44]. Due to the bilateral symmetry of the animal and human body [27, 67], i.e. right and left sides are mirror shapes of one another, the BSE can be defined as a non-trivial chiral shape matching problem. Then, given a known reference or source bone, assessing the side of a target bone involves finding the best match between the source-target pair and its mirrored counterpart. The complexity of this task arises from subtle local differences in mirrored bones, which makes it a suitable application of the proposed shape matching technique utilizing coupled Laplacian eigenmaps as local geometry descriptors. We propose a benchmark for human BSE by extracting surface point clouds of different bones, e.g. femur, hip and tibia, from public datasets [18, 31, 47]. We also discuss, in the Supplementary Material, a non-rigid correspondence application of the proposed matching algorithm, the cross-species BSE (human to animal), on an internal dataset [40].
@@ 2 Related Works
Early efforts in developing hand-crafted 3D local features for point cloud matching, registration and retrieval have typically drawn inspiration from 2D descriptors. Several methods, such as Signature of Histograms of OrienTations (SHOT) [58, 66], Rotational Projection Statistics (RoSP) [25] and Unique Shape Context (USC) [66] rely on the estimation of an unique Local Reference Frame (LRF). The latter is usually based on the eigendecomposition of the covariance matrix of the neighbours of a point of interest, which are then projected into the LRF to analyze their geometric properties. For instance, SHOT [58] captures local shape and orientation information by computing histograms of surface normals and point distribution around keypoints. In contrast, LRF-free approaches [10, 56, 57] try to rely just on features that are intrinsically invariant. The most common LRF-free features are Fast Point Feature Hisograms (FPFH) [57] which, similarly to SHOT, calculates histograms of surface normals considering their relationships in a local region around a keypoint and generating $33$-dimensional descriptors. Despite the progress made with hand-crafted 3D local features, they encounter difficulties when it comes to deal with issues like point cloud resolutions, noisy data, and occlusions [26].
@@
On the contrary, spectral-based point cloud descriptors [1, 28, 52, 70, 62, 55, 69] are a category of feature extraction methods that leverage spectral analysis techniques from graph theory to capture the underlying structure and intrinsic geometric properties of point clouds. Among them, shape-DNA [52] is a surface descriptor based on eigenvalue analysis of the Laplace-Beltrami operator. They propose to use the sequence of eigenvalues (spectrum) of the Laplace operator as a fingerprint characterizing the intrinsic geometry of 3D shapes represented as point clouds. This method has been used for shape retrieval, classification, and correspondence. Weinmann et al. [70] also proposed to extract a features set consisting of $8$ eigenvalues-based indices for each 3D point of a cloud. Furthermore, Heat Kernel Signature (HKS) [62] and Wave Kernel Signatures (WKS) [1] are descriptors measuring how heat and wave propagate across a shape, having the eigendecomposition as leading element of the computation. Scaled eigenvectors evaluated at each point are instead directly exploited by the Global Point Signature (GPS) [55] to represent a point cloud as a set of infinite-dimensional vectors, characterizing each point within the global context of the surface it belongs. Nevertheless, the vast majority of these works assumes that the eigenvalues of a shape are distinct and, therefore, can be ordered. Indeed, in practice, due to numerical approximations, we cannot guarantee that the eigenvalues of the Laplacian are all distinct and, possible symmetries in the shapes may cause some of them to have multiplicity grater than one. As shown by Mateus et al. [42], when dealing with shape matching, an elegant way to overcome this problem is to use the Laplacian eigenmaps scheme [4], which can be seen as a reduced GPS, and perform a-posteriori alignment of the resulting point cloud embeddings. Matching the eigenfunctions histograms, i.e. their signatures, is the only reliable method for such embeddings alignment [42, 60]. Recently, Ma et. al. [39] proposed a canonization algorithm for sign and basis invariance, called Maximal Axis Projection (MAP), that adopts the permutation-invariant axis projection functions to determine the canonical directions. Unfortunately, when the eigenvalues separation is small due to symmetries in the geometries, these methods becomes unstable and sensitive to noise. To overcome this issue, in this work we propose the Coupled Laplacian operator to produce a-priori aligned eigenmaps for several registered shapes, and perform locally-aware shape matching.
@@
Deep-learned point-cloud descriptors [24, 72, 6, 77] arose as an alternative to generate local features for 3D sufraces. We can distinguish three main categories depending on the backbone architecture employed. Convolutional Neural Networks (CNNs) are often used on point clouds projected into 2D depth images [16, 29] or directly on 3D voxels [24, 77, 29]. Secondly, to work directly on raw point cloud data, PointNet and PointNet++ have been proposed aiming to learn rotation and permutation invariant features [49, 50]. Based on PointNet, several learned descriptors have been introduced [14, 13, 75, 76]. Among them, PPFNet [14] and PPF-FoldNet [13] try to improve the feature representation of PointNet by incorporating global context and point-pair features. Nevertheless, the lack of convolutional layers in these models limits the learning of local geometries. Finally, transformers-based descriptors, such as Deep Closest Point [68], have been recently proposed trying to exploit the attention mechanisms to capture shapes and surfaces intrinsic characteristics. Arguably, training a deep-learning model, especially if based on transformers, it is not always feasible in terms of training samples required to achieve good performances. Furthermore, in supervised algorithms [77, 24, 14, 13], paired 3D patches, such as in the 3DMatch dataset [77], are needed to train the models. For this reason, the 3D rigid point cloud matching method we propose does not need training and utilizes only the similarity among properly aligned spectral embeddings to relate an unseen target shape to the given references.
@@ 3 Method
where $\boldsymbol{B}\in\mathbb{R}^{n\times n}$ is generally set as $\boldsymbol{D}$, and $\lambda_{i}\leq\lambda_{i+1}$. Finally, eigenmaps are simply eigenvectors sub-spaces, generated by leaving out $\boldsymbol{\phi}_{0}$, corresponding to $\lambda_{0}=0$, and using the next $m$ eigenvectors for embedding graph nodes in an $m$-dimensional space, $\boldsymbol{x}_{i}\rightarrow[\boldsymbol{\phi}_{1}(i),\cdots,\boldsymbol{\phi}_{m}(i)]$
@@
Coupled Laplacian. The complete pipeline to generate aligned Laplacian embeddings through graph coupling is illustrated in Fig. 2. Let $\mathcal{G}^{\mathcal{T}}$ be the graph of a target 3D point cloud and $\mathcal{G}^{S_{1}},\cdots,\mathcal{G}^{S_{N}}$ the ones of $N$ different sources. We construct a global graph, or coupled graph, $\mathcal{G}^{C}$, by adding cross-connections separately between the vertices of the source, $\mathcal{V}^{\mathcal{T}}=\{\boldsymbol{x}^{\mathcal{T}}_{i}\}_{i=1}^{n^{\mathcal{T}}}$, and each of the reference shapes, $\mathcal{V}^{S_{k}}=\{\boldsymbol{x}^{S_{k}}_{i}\}_{i=1}^{n^{S_{k}}}$ with $k=1,\cdots,N$. To include meaningful cross-connections, we first perform a rigid [19, 20, 45] or affine [45] registration of the source geometries to align to the target. These methods are preferred to a non-rigid registration because, when dealing with the identification of fine local variations, they do not change the relative position of the points inside a point-set and, therefore, the deformations are kept unchanged. After that, a sub-set of vertices, $F^{\mathcal{T}}\subset\mathcal{V}^{\mathcal{T}}$ , is stochastically extracted from the target geometry and their nearest correspondences are searched in each of the aligned reference point clouds as
$F^{S_{k}}=\{\boldsymbol{f}^{S_{k}}_{j}:\operatorname*{arg\,min}_{\boldsymbol{x}^{S_{k}}_{i}\in\mathcal{V}^{S_{k}}}\text{d}^{2}(\boldsymbol{x}^{S_{k}}_{i},\boldsymbol{f}^{\mathcal{T}}_{j})|\boldsymbol{f}^{\mathcal{T}}_{j}\in F^{\mathcal{T}}\}$ (3)
@@
with $k=1,\cdots,N$. In this way, there is no constraint on the original number points comprising each shape nor on their initial coordinates systems. Furthermore, the cardinality of the target sub-set, is chosen as a fraction of its total points, $|F^{\mathcal{T}}|=l\cdot n^{\mathcal{T}}$, where $0<l\leq 1$. The cross-connection are therefore added for each pair $(\boldsymbol{f}^{S_{k}}_{i},\boldsymbol{f}^{\mathcal{T}}_{i})$, $1\leq i\leq|F^{\mathcal{T}}|$, between the target and the $k$-th source. Finally, we arrange the vertex indices of the coupled graph such that they are grouped for each individual geometry, and the whole weighted adjacency matrix can be computed as in Eq. 1.
@@
We define the Laplacian matrix of a global graph, constructed as describe above, as the Coupled Laplacian, $\boldsymbol{L}^{C}\in\mathbb{R}^{n\times n}$, where $n=n^{\mathcal{T}}+\sum_{k=1}^{N}n^{S_{k}}$. Thanks to the vertex ordering of the coupled graph, the solution of the generalized eigenproblem of Eq. 2 applied to the Coupled Laplacian, with $\boldsymbol{B}^{C}=\text{diag}([\boldsymbol{D}^{\mathcal{T}},\boldsymbol{D}^{S_{1}},\cdots,\boldsymbol{D}^{S_{N}}])$, yields to coupled eigenvectors $\{\boldsymbol{\phi}^{C}_{i}\}_{i=0}^{n}$. These eigenvectors can then be split into each single component of the coupled graph as
@@
Shape Matching with Eigenmaps. The Coupled Laplacian allows the generation of $m$-dimensional embeddings aligned for the graph vertices of the target and reference shapes, $\boldsymbol{\Phi}^{\mathcal{T}}=(\boldsymbol{\phi}^{\mathcal{T}}_{1},\cdots,\boldsymbol{\phi}^{\mathcal{T}}_{m})\in\mathbb{R}^{n_{T}\times m}$ and $\boldsymbol{\Phi}^{S_{k}}=(\boldsymbol{\phi}^{S_{k}}_{1},\cdots,\boldsymbol{\phi}^{S_{k}}_{m})\in\mathbb{R}^{n_{S_{k}}\times m}$ for $1\leq k\leq N$, respectively. Therefore, a comparison of multiple geometries in this higher dimensional spectral space yields to a proper consideration of local structures. We consider in the following two possible applications of the coupled eigenmaps: (1) a global shape matching score and (2) local similarity scores. The first one can be obtained by measuring the similarity between the aligned eigenspaces through the Grassmann distance, $d_{G}(\cdot,\cdot)$ [74]. Nevertheless, it cannot be computed directly, since we consider point clouds of arbitrary size and we do not have point-to-point correspondences. Hence, we restrict the distance computation only to the set of cross-connected vertices, $F^{\mathcal{T}}$ for the target and $F^{S_{k}}$ for the $k$-th reference. In order to get the reduced basis restricted only to those coupled points, we perform a QR decomposition of the restriction of the eigenmodes to the cross-points. The best matching source is then given by
@@
where $\boldsymbol{Q}^{\mathcal{T}}$ and $\boldsymbol{Q}^{S_{k}}$ are obtained from the QR factorization of $\boldsymbol{\Phi}^{\mathcal{T}}(F^{\mathcal{T}},:)$ and $\boldsymbol{\Phi}^{S_{k}}(F^{S_{k}},:)$, respectively. This first approach is used in the following to solve the BSE task by using two source shapes, i.e. the reference bone of known side and its contralateral mirrored version.